OpenStudy (anonymous):

find the first, fourth, and eighth terms of the sequence A(n) = -5*3^n-1

OpenStudy (helder_edwin):

u have \[\large A(n)=-5\cdot3^{n-1} \] for the first term set n=1 and u get \[\large A(1)=-5\cdot3^{1-1}=-5\cdot3^0=-5\cdot1=-5 \]

OpenStudy (helder_edwin):

now. for the 4th term set n=4 and for the 8th term set n=8.

OpenStudy (anonymous):

these are the answers: a) -15; -405;-32.805; b) 1;-3,375; -170,895,375 c) -5; -135; -10,935 d) -45; -1,215; -98,415